# RECAST: Strengthening LLMs’ Complex Instruction Following with Constraint-Verifiable Data

### v2

@@ C.4 Details of SFT
This section details our model training configuration based on the LlamaFactory [50] framework. We employed Supervised Fine-Tuning (SFT) with a maximum sequence length of 4096 tokens. For optimization, we used a linear learning rate scheduler with a peak learning rate of 2.0e-5, 3% warmup ratio, and trained for 3 epochs. We configured training with a per-device batch size of 4 and gradient accumulation steps of 32, resulting in an effective batch size of 128. To optimize training efficiency, we utilized DeepSpeed with ZeRO stage 3 optimization and bfloat16 mixed precision. All experiments were conducted on a cluster of 8 NVIDIA H800 GPUs.
@@ C.5 Details of RLVC
For reinforcement learning, we implemented our RLVC approach based on the VeRL [51] framework with Group Relative Policy Optimization (GRPO). Our configuration used a learning rate of 1e-6 with maximum sequence lengths of 1024 tokens for both prompts and responses. We employed a batch size of 512 with PPO mini-batches of 128 and micro-batches of 16 per GPU. For optimization stability, we incorporated KL divergence regularization with a coefficient of 0.001 using the low-variance KL implementation, while enabling gradient checkpointing for memory efficiency. The rollout process utilized tensor model parallelism with a size of 2 and vLLM for acceleration, generating 16 candidate responses per prompt with 60% GPU memory utilization. We integrated our custom constraint-specific reward functions through the thread-based reward manager with asynchronous reward calculation. All experiments were conducted on 8 GPUs for 200 optimization steps.

### v3

@@ C.4 Details of SFT
This section details our model training configuration based on the LlamaFactory Zheng et al. [2024] framework. We employed Supervised Fine-Tuning (SFT) with a maximum sequence length of 4096 tokens. For optimization, we used a linear learning rate scheduler with a peak learning rate of 2.0e-5, 3% warmup ratio, and trained for 3 epochs. We configured training with a per-device batch size of 4 and gradient accumulation steps of 32, resulting in an effective batch size of 128. To optimize training efficiency, we utilized DeepSpeed with ZeRO stage 3 optimization and bfloat16 mixed precision. All experiments were conducted on a cluster of 8 NVIDIA H800 GPUs.
@@ C.5 Details of RLVC
For reinforcement learning, we implemented our RLVC approach based on the VeRL Sheng et al. [2024] framework with Group Relative Policy Optimization (GRPO). Our configuration used a learning rate of 1e-6 with maximum sequence lengths of 1024 tokens for both prompts and responses. We employed a batch size of 512 with PPO mini-batches of 128 and micro-batches of 16 per GPU. For optimization stability, we incorporated KL divergence regularization with a coefficient of 0.001 using the low-variance KL implementation, while enabling gradient checkpointing for memory efficiency. The rollout process utilized tensor model parallelism with a size of 2 and vLLM for acceleration, generating 16 candidate responses per prompt with 60% GPU memory utilization. We integrated our custom constraint-specific reward functions through the thread-based reward manager with asynchronous reward calculation. All experiments were conducted on 8 GPUs for 200 optimization steps.